Let us consider a quadratic polynomial f(x) such that equation f(x) = 6x - 18 has exactly one
root, and equation f(x) = 9 - 3x has exactly one root. Find the minimum value of a
discriminant of f(x).
1
Expert's answer
2020-11-09T20:12:36-0500

Let "f(x)=ax^2+bx+c, a\not=0." Then the discriminant if "f(x)" is


"D=b^2-4ac"

The equation "f(x)=6x-18" has exactly one root


"ax^2+bx+c=6x-18"

"ax^2+(b-6)x+(c+18)=0"

"(b-6)^2-4a(c+18)=0"

The equation "f(x)=9-3x" has exactly one root


"ax^2+bx+c=9-3x"

"ax^2+(b+3)x+(c-9)=0"

"(b+3)^2-4a(c-9)=0"


We have the system


"b^2-12b+36-4ac-72a=0""b^2+6b+9-4ac+36a=0"

"b^2-4ac=12b+72a-36""b^2-4ac=-6b-36a-9"

"12b+72a-36=-6b-36a-9"

"b=-6a+1.5""b^2-4ac=36a-9-36a-9"

"D=-18"



The value of the discriminant of "f(x)" is "-18."